OpenStudy (anonymous):

negative 6 over 5, multiplied by x minus 4 equals negative 46

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

so -6/5 * x - 4 = -46

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look about right

OpenStudy (anonymous):

so how do I solve for x

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

\[\frac{ -6 }{ 5 }x -4=-46\]

OpenStudy (anonymous):

Step 1: Add 4 to both sides.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tell me what you get when you do that step

OpenStudy (anonymous):

-50

OpenStudy (anonymous):

no -42

OpenStudy (anonymous):

no adding negatives makes the negative go up

OpenStudy (anonymous):

yes when you ad 2 negatives but this is a negative and a positive

OpenStudy (anonymous):

does that make sense @KatelynisAM4ZING